OpenStudy (anonymous):

let f(x)=x/x+4 solve the quadratic type equation f(x)-13 square root of f(x) +36=0 I have no idea where to start

OpenStudy (anonymous):

\[Let f(x)=\frac{ x }{ x+4 }. Solve the quadratic type equation f(x)-13\sqrt{f(x)}+36=0\]